OpenStudy (anonymous):

use identites to find the exact value of the indicated trigonometric function of the acute angle delta. sin delta= 2squr.2 over 3, cos delta=1 over 3 find csc delta

OpenStudy (anonymous):

\[\sin \theta=\frac{ 2\sqrt{2} }{ 3 } ,\cos \Theta=\frac{ 1 }{ 3}\] find csc